OpenStudy (anonymous):

Sandy claims that 50!+31 is a prime number because none of the divisibility tests work and to her it looks prime. Sandy is wrong. Explain why 50!+31 is not prime and why her explanation is wrong

OpenStudy (anonymous):

prime number should be the one (50! + 3!) + 1

OpenStudy (anonymous):

no.. it goes like this...

OpenStudy (anonymous):

\[50!=50\times49\times48\times\ldots\times32\times31\times30\times\ldots\times2\times1\] so, 31 is a factor of 50! so, let \(50! = 31\times c\) \[(50!+31)=(31\times c+31)=31(c+1)\] this number is clearly divisible by 31 and hence not a prime

OpenStudy (anonymous):

Inkyvoyd, I have some additional questions. Think you could help me out?